OpenStudy (anonymous):

find the sets of values of "x" for which 2x^2-4x-3 is greater than "x"

OpenStudy (anonymous):

x < -1/2 or x > 3

OpenStudy (anonymous):

that is x belongs to the set (-1/2,3)

OpenStudy (sriram):

We want 2x^-4x-3 -x >0 hence f(x)= 2x^2-5x-3 f(x)=0 at x=-1/2 &3 f'(x)=4x-5 this is +ve for x>5/4 so for x>5/4 the quadratic is greater than x